OpenStudy (anonymous):

What is the continuity test? I need to know how to solve this: given g(x) = {2x^3 -1 , x < 1/3 {bx^2 +1 , x> or equal to 1/3 find b such that g(x) is a continuous function. this is bc calc.

OpenStudy (zzr0ck3r):

when we plug in 1/3 to the top we get 2/27-1 = -25/27 we want the same thing to happen on the second part b(1/3)^2+1 = -25/27 solve for b

OpenStudy (anonymous):

but why the same thing on the second part?

OpenStudy (zzr0ck3r):

well, we want it to be continuous and 2x^3-1 is continuous and so is the bottom one for any number b but we dont want there to be a jump when we go from the first function defined for x<1/3 to the second function. so we make sure they have the same value at the point we have a break at.

OpenStudy (zzr0ck3r):

go to wolfram and graph them both and use different values for b, I think you will understand it better

OpenStudy (zzr0ck3r):

this is just an example but say the graph looked like these  the one on the left is the top and the one on the right is the bottom for some value b now if we find the right b we will have

OpenStudy (zzr0ck3r):